# Diffusion-Augmented Contrastive Learning: A Noise-Robust Encoder for Biosignal Representations

### v1

@@ III-D Mathematical Formulation
The central component of the learning phase is a U-Net encoder, $\text{Enc}_{\theta}$, designed to take a noisy latent vector $\mathbf{z}_{t}$ and its corresponding timestep $t$ as input, allowing the network to condition its output on the noise level. For a given triplet, views are generated as follows: an anchor and a positive view are created from the same initial latent vector ($\mathbf{z}_{0,A}=\mathbf{z}_{0,P}$) by applying the diffusion process with independently sampled noise and timesteps ($t_{A},t_{P}$). A negative view is generated from a latent vector of a different class, $\mathbf{z}_{0,N}$, at a random timestep $t_{N}$. These three noisy views ($\mathbf{z}_{t_{A},A}$, $\mathbf{z}_{t_{P},P}$, $\mathbf{z}_{t_{N},N}$) are then passed through the shared U-Net encoder to produce the final embeddings $\mathbf{e}_{A},\mathbf{e}_{P},\mathbf{e}_{N}$. The encoder is trained by minimizing a supervised triplet margin loss:
@@ III-E Training and Inference Procedure
  • Inference: For a given test sample $\mathbf{x}_{test}$, it is passed through the frozen VAE encoder to obtain the corresponding clean latent vector $\mathbf{z}_{0,test}$. This vector is then passed through the trained U-Net encoder using a fixed, low-noise timestep (e.g., $t=1$) to produce a stable, noise-robust final embedding $\mathbf{e}_{test}$. A normal-class prototype, $\boldsymbol{\mu}_{normal}$, is pre-computed by taking the mean of the final embeddings of all normal samples from the training set. The final anomaly score is determined by the Euclidean distance $\|\mathbf{e}_{test}-\boldsymbol{\mu}_{normal}\|_{2}$.
@@ VI Conclusion
This work introduced Diffusion-Augmented Contrastive Learning (DACL), a novel framework that successfully synthesizes diffusion models and supervised contrastive learning. We proposed using the diffusion forward process as a principled and learned data augmentation technique to generate the ”views” required for a contrastive learning objective. Our assessment of the PhysioNet 2017 dataset confirmed the efficacy of our method, attaining a competitive AUROC of 0.7815. This methodological study, although not surpassing fully supervised approaches, successfully developed a novel paradigm for learning noise-invariant representations that underpin class discrimination. This study primarily contributes conceptually, paving the way for further research in noise-robust representation learning for biosignals and other fields where augmentations are complex. Subsequent research should focus on exploring more sophisticated diffusion schedules, applying the framework to other challenging time-series domains, and investigating adaptive timestep sampling strategies during training.

### v2

@@ III-D Mathematical Formulation
The central component of the learning phase is a U-Net style encoder, $\text{Enc}_{\theta}$, implemented here as a Multi-Layer Perceptron (MLP) suitable for operating on vector-based latent spaces. While not structurally a convolutional U-Net, we adopt this terminology to reflect its functional role as the conditional noise predictor, which is analogous to the U-Net’s role in seminal diffusion models. For a given triplet, views are generated as follows: an anchor and a positive view are created from the same initial latent vector ($\mathbf{z}_{0,A}=\mathbf{z}_{0,P}$) by applying the diffusion process with independently sampled noise and timesteps ($t_{A},t_{P}$). A negative view is generated from a latent vector of a different class, $\mathbf{z}_{0,N}$, at a random timestep $t_{N}$. These three noisy views ($\mathbf{z}_{t_{A},A}$, $\mathbf{z}_{t_{P},P}$, $\mathbf{z}_{t_{N},N}$) are then passed through the shared U-Net encoder to produce the final embeddings $\mathbf{e}_{A},\mathbf{e}_{P},\mathbf{e}_{N}$. The encoder is trained by minimizing a supervised triplet margin loss:
@@ III-E Training and Inference Procedure
  • Inference: For a given test sample $\mathbf{x}_{test}$, it is passed through the frozen VAE encoder to obtain the corresponding clean latent vector $\mathbf{z}_{0,test}$. This vector is then passed through the trained U-Net encoder using a fixed, low-noise timestep (e.g., $t=1$) to produce a stable, noise-robust final embedding $\mathbf{e}_{test}$. A normal-class prototype, $\boldsymbol{\mu}_{normal}$, is pre-computed by taking the mean of the final embeddings of all normal samples from the training set. The anomaly score is determined by the Euclidean distance $\|\mathbf{e}_{test}-\boldsymbol{\mu}_{normal}\|_{2}$.
@@ VI Conclusion
This work introduced Diffusion-Augmented Contrastive Learning (DACL), a novel framework that successfully synthesizes diffusion models and supervised contrastive learning. We proposed using the diffusion forward process as a principled and learned data augmentation technique to generate the ”views” required for a contrastive learning objective. Our assessment of the PhysioNet 2017 dataset confirmed the efficacy of our method, attaining a competitive AUROC of 0.7815. This methodological study, although not surpassing fully supervised approaches, successfully developed a novel paradigm for learning noise-invariant representations that underpin class discrimination. This study aims to pave the way for further research in noise robust representation learning for biosignals and other fields where augmentations are complex.